OpenStudy (anonymous):

Using complete sentences, explain how to find the zeros of the function f(x) = 2x^3-9x+3.

OpenStudy (anonymous):

well I think you factor out the equation and use the ( ? ) ( ? ) to solve for x, whatever the x's are will be your zeros

OpenStudy (anonymous):

what is the (?)(?)

OpenStudy (anonymous):

factorize first

OpenStudy (anonymous):

how

hero (hero):

Start by finding two numbers, a and c, such that: a + c = -9 ac = (2)(3)

hero (hero):

Or you can just graph the function and observe where it crosses the x-axis

hero (hero):

Yet another alternative is to complete the square. If you want to do that, you can start by setting f(x) to 0.

hero (hero):

2x^2 - 9x + 3 = 0

hero (hero):

One of the above are methods students use to find the zeroes.

hero (hero):

Okay, that's a good start.

OpenStudy (anonymous):

i cant seem to factor the 2x^2 - 9x + 3 = 0

hero (hero):

That's because it is not factorable by methods you normally use.

hero (hero):

If the quadratic function isn't factorable, you usually use the quadratic formula to find the zeros.

hero (hero):

\(\large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\)

OpenStudy (anonymous):

im doing that right now

hero (hero):

Great :D

OpenStudy (anonymous):

\[9+√57/4 \]

OpenStudy (anonymous):

\[9-√57/4\]

hero (hero):

And that's still simplifyable

OpenStudy (anonymous):

ohh

hero (hero):

Did you figure out your mistake while using the formula?

OpenStudy (anonymous):

nope

hero (hero):

looks like I'm the one that made the mistake.

hero (hero):

Good job

OpenStudy (anonymous):

hahaha thanks!